# Probing Mechanical Reasoning in Large Vision Language Models

### v2

@@ A.2 Pulley Systems
VLMs perform poorly in recognizing movable pulley systems. In Experiment 8A, the image includes a standard single movable pulley system and a standard single fixed pulley system. The question "Which system requires less effort?" is essentially asking whether the model can correctly select the movable pulley. Clearly, the model failed in its selection, as it straightforwardly provided an incorrect answer in its explanation. VLMs also struggle in predicting whether a suspended weight is being lifted or lowered through a pulley system. Experiments 8B-F either directly or indirectly reflect this issue, with Experiment 8D being the most direct and concise. In Experiment D, the weight is directly attached to the movable pulley, and by pulling the other end of the rope, the pulley and the weight are lifted. However, the model’s response was the exact opposite of the correct answer. In its explanation, the model seemed to imply that the pulley was not fixed (though it did not explicitly state that it was a movable pulley), and the physics it provided was entirely incorrect. Therefore, we can hypothesize that the model’s poor performance in predicting the weight’s movement may be due to its limited ability to recognize movable pulleys. However, the specific reasons require further experiments to be analyzed in detail.
@@ A.5 Inertia and Motion
Experiment 11E involves scene-based question. The model not only identified that the cart was not stationary, but it also further analyzed that if the object suspended on the cart were stationary, it would fall vertically due to gravity. However, the depicted scene likely indicates that the cart is accelerating forward. Experiments 11A, 11B, and 11D, on the other hand, are cases of simple extrapolation. They are categorized as "simple extrapolation" because the scenarios are straightforward and involve a single, clear change. For example, in Experiment 11A, the change involves the cart suddenly stopping—a relatively simple scene (involving only two objects, the drink and the cart, with straight-line motion and vertical force equilibrium). Similar patterns apply to Experiments 11B and 11D.

### v3

@@ A.2 Pulley Systems
VLMs perform poorly in recognizing movable pulley systems. In Experiment 8A, the image includes a standard single movable pulley system and a standard single fixed pulley system. The question ”Which system requires less effort?” is essentially asking whether the model can correctly select the movable pulley. Clearly, the model failed in its selection, as it straightforwardly provided an incorrect answer in its explanation. VLMs also struggle in predicting whether a suspended weight is being lifted or lowered through a pulley system. Experiments 8B-F either directly or indirectly reflect this issue, with Experiment 8D being the most direct and concise. In Experiment D, the weight is directly attached to the movable pulley, and by pulling the other end of the rope, the pulley and the weight are lifted. However, the model’s response was the exact opposite of the correct answer. In its explanation, the model seemed to imply that the pulley was not fixed (though it did not explicitly state that it was a movable pulley), and the physics it provided was entirely incorrect. Therefore, we can hypothesize that the model’s poor performance in predicting the weight’s movement may be due to its limited ability to recognize movable pulleys. However, the specific reasons require further experiments to be analyzed in detail.
@@ A.5 Inertia and Motion
Experiment 11E involves scene-based question. The model not only identified that the cart was not stationary, but it also further analyzed that if the object suspended on the cart were stationary, it would fall vertically due to gravity. However, the depicted scene likely indicates that the cart is accelerating forward. Experiments 11A, 11B, and 11D, on the other hand, are cases of simple extrapolation. They are categorized as ”simple extrapolation” because the scenarios are straightforward and involve a single, clear change. For example, in Experiment 11A, the change involves the cart suddenly stopping—a relatively simple scene (involving only two objects, the drink and the cart, with straight-line motion and vertical force equilibrium). Similar patterns apply to Experiments 11B and 11D.